OpenStudy (anonymous):

Solve the equation, find an explicit solution if possible. Equation: x(dy/dx)-x^3e^(2x)-2y=0

OpenStudy (jamesj):

If you don't know how to approach this problem, watch this great lecture. It will explain to you exactly how to solve this sort of problem. http://ocw.mit.edu/courses/mathematics/18-03-differential-equations-spring-2010/video-lectures/lecture-3-solving-first-order-linear-odes/

zepdrix (zepdrix):

\[\large xy'-x^3e^{2x}-2y=0\]Move some stuff around,\[\large xy'-2y=x^3e^{2x}\]Divide both sides by x to get it off of the leading y term.\[\large y'+\left(-\frac{2}{x}\right)y=x^2e^{2x}\]

zepdrix (zepdrix):

From here we need to find an integrating factor:\[\large \mu=e^{-\int\limits\frac{2}{x}dx}\]

OpenStudy (jamesj):

Stop there please. It is important that the questioner have the chance to take the next steps themselves.

OpenStudy (anonymous):

ok so e=1/x^2 right?

OpenStudy (anonymous):

sorry, μ=e^(ln(x^2))

OpenStudy (anonymous):

then μ=1/x^2 and then you multiply it on both sides, right?

OpenStudy (jamesj):

Yes

OpenStudy (anonymous):

you get \[(y'/x^2)-(2y/x^3)=e^(2x)\]

OpenStudy (anonymous):

then i'm lost...

zepdrix (zepdrix):

Hmm looks good so far! :) The point of introducing an `integrating factor` is that it SHOULD turn our left side into something that looks like the product rule. We're basically trying to apply the product rule in reverse. We have something like this:\[\large u'v+uv'\] and we want to try it like this,\[\large (uv)'\]

zepdrix (zepdrix):

want to write it like** blah typo :3

OpenStudy (jamesj):

What's the whole point of the integrating factor? ... right, what zepdrix wrote. I strongly recommend you watch the lecture where this is carefully and nicely explained.

zepdrix (zepdrix):

If you correctly found your integrating factor \(\mu\), then the left side should simplify down to,\[\large \left(\mu \cdot y\right)'\]It's a good idea to apply the product rule from here, to make sure it gives you the two terms you had. That will let you know if you made a mistake! :)

OpenStudy (anonymous):

ok I'm going to substitute -2/x^3 with d/dx (1/x^2) which will look like \[y* d/dx (1/x^2)+(dy/dx)/x^2=e^(2x)\]

OpenStudy (anonymous):

I get d/dx (y/x^2)=e^(2x) which i integrate both sides and get y/(x^2) = e^(2x)/2+\[c_{1}\]

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

and then i can divide both sides by 1/x^2

OpenStudy (anonymous):

woot thanks!

zepdrix (zepdrix):

yay good job c: